Solve y"' -5y"+ 4y'=8ex + 4x.



1
Expert's answer
2021-07-19T05:46:22-0400

Let "y'=z." Then


"z''-5z'+4z=8e^x+4x"

The homogeneous equation is


"z''-5z'+4z=0"


The characteristic equation is


"r^2-5r+4=0"

"r_1=1, r_2=4"

The general solution of the homogeneous differential equation is


"z_h=C_1e^x+C_2e^{4x}"

Find the particular solution of the nonhomogeneous equation in form


"z_p=Axe^x+Bx+C"

Then


"z_p'=Ae^x+Axe^x+B"

"z_p''=2Ae^x+Axe^x"

Substitute


"2Ae^x+Axe^x-5Ae^x-5Axe^x-5B"

"+4xAe^x+4Bx+4C=8e^x+4x"


"xe^x:0=0"

"e^x:-3A=8"

"x^1:4B=4"

"x^0:-5B+4C=0"

"z_p=-\dfrac{8}{3}xe^x+x+\dfrac{5}{4}"

The general solution of the differential equation


"z''-5z'+4z=8e^x+4x"

is


"z=C_1e^x+C_2e^{4x}-\dfrac{8}{3}xe^x+x+\dfrac{5}{4}"

Integrate


"y=\int(C_1e^x+C_2e^{4x}-\dfrac{8}{3}xe^x+x+\dfrac{5}{4})dx"

"\int xe^xdx=xe^x-e^x+C_3"

"y=c_1+c_2e^x+c_3e^{4x}-\dfrac{8}{3}xe^x+\dfrac{1}{2}x^2+\dfrac{5}{4}x"